When you go out to your car one cold winter morning you discover a 0.60-cm thick layer of ice on the windshield, which has an area of 1.3 m2 . If the temperature of the ice is -6.0 ∘C, and its density is 917 kg/m3, find the heat required to melt all the ice.

Express your answer using two significant figures.
Answer #1

Thickness of layer, x =0.6 cm ,

Area A =1.3 m^2

Volume,

V= A.x =1.3*0.006 = 0.0078 m^3

densityof ice , d =917 kg/m^3

Calculate mass of ice

m =d*V = 917*0.0078

= 7.1526 kg

T1 =-6 oC, T2 = 0 oC
specific heat of ice Cp = 2108 J/kg.K

Latent heat of ice L = 333 kJ/kg

Calculate  the heat required to melt all the ice.

Q =mCpdT +mL

Q = [ 7.1526*2108*(0+6) ] + (7.1526*333*1000)

Q = 2.5 x106 J